# Conceptualization, Data Curation

### v2

@@ 5.2 Unseen Data
Deep learning-based FDD algorithms can not perform well on unseen operation modes and faults. To deal with previously unknown faults, an algorithm based on supervised contrastive learning (SCL) is used in [66], in which the normal samples are contrasted with negative augmentations of themselves. Contrastive Learning is a deep learning technique for unsupervised representation learning. The goal is to learn a representation of data such that similar instances are close together in the representation space, while dissimilar instances are far apart. Soft Brownian Offset (SBO), sampling and shrinkage autoencoder are used to generate the out-of-distribution (OOD) samples, to obtain some information from unknown classes. The goal is to prevent unseen fault samples from being classified as normal conditions. Balanced fault diagnosis, imbalanced fault diagnosis, and few-shot diagnosis of TEP are tested by the proposed method, demonstrating good performances. Faults 1, 2, 4, 6, and 7 are considered as the seen faults, and faults 8, 12, 14, and 18 are selected as the unseen faults. The detection accuracies in the balanced scenario are 97.5%, 100%, 100%, and 90.63%. In the imbalanced scenario, where the balance ratio of normal to fault is 10:1, the accuracies are 97.2%, 98.75%, 100%, 90.38%, In few-shot learning, where there is only 15 samples of each fault, the accuracies are 95.1%, 97.6%, 99.86%, 88.2% respectively. Xiao et al. [98] implement a developed domain generalization (DG) method that guarantees generalization over unseen domains. DG learns a universal fault diagnosis model from historical operating modes and can generalize to unseen modes. A novel labelling and class progressive adversarial learning (LCPAL) algorithm for FDD is introduced to learn domain-invariant feature representations. This model aligns the label distribution by weighting the source classification loss based on the number of instances in each category. Stacked autoencoders can be used for handling incomplete, missing or unlabeled data. In [27] a modified stacked autoencoder is used to learn low-feature representations from incomplete data input. Only 40 input variables out of 52 are randomly chosen for fault detection and the average accuracy for the fault diagnosis of faults 1, 4, 7, 8, 10, 13 is 83.76%.
5.3 Imbalanced Data
@@
Industrial systems usually work in normal condition, and normal data is ample. However, faults occur infrequently and the number of fault samples is rarely equal to the number of normal samples. AI algorithms can not perform well when the data is imbalanced. The performance of the data-driven FDD algorithms depends on the availability of the data, however, in many real-world cases, data is imbalanced or incomplete. To deal with this problem and achieve good FDD accuracy, different approaches are introduced. One approach for handling imbalanced data sets is generating synthetic samples of the minor class to make the data balanced [3]. The other approach is using feature extraction methods focusing on available data and considering one-class classification and anomaly detection methods.
@@
Tian et al.[86], used a generative adversarial network (GAN) to reconstruct missing data, Spearman’s rank correlation coefficient (SRCC) for dimension reduction, select the most important features and noise cancelling, and finally DBN for FDD. The average FDR of the process for 20 faults of TEP is 89.70% which is 6.23% higher than the traditional DBN method. Hu et al.[32], proposed an incremental imbalance modified deep neural network (incremental-IMDNN) in combination with active learning for FDD of imbalanced data. An imbalance modified synthetic minority oversampling technique (IM-SMOTE) is used for addressing data imbalance. The model extends the structure hierarchically with the arrival of new faults. Similar faults are merged into the same classes by using fuzzy clustering. To tackle the class imbalance problem in FDD of chemical processes, an Imbalance Modified Convolutional Neural Network combined with incremental learning, as described in [26], generates new samples from imbalanced raw data using a dynamic modified synthetic minority oversampling technique (DM-SMOTE). This model updates itself by arriving the new data.
@@
To identify unknown faults a few-shot learning-based unknown recognition and classification (FSLB-UR&C) approach is introduced in [57]. For feature extraction and comparing the similarity of features, a gated recurrent unit with a dot-product attention layer is used. This method does not require a large amount of data for training. The classification accuracy of the known faults of TEP is 88.9%, and for unknown faults is 85.7%. The known faults include Faults 0, 2, 12, and 14, while the unkown faults include Faults 1, 4, 5, 7, 8, 10, 11, and 13. Xu et al. [101], propose a few-shot learning fault diagnosis method based on a class-rebalance strategy, to diagnose new scarce faults. The fault samples are transferred into a feature space, where each type of fault, even each new instance, is mapped into a separate feature cluster. In this way, fault diagnosis can be achieved by estimating feature similarity between instances and faults. 1D-CNN with an attention module is used for feature extraction. Data augmentation and class rebalance are done to increase the fault samples of scarce classes. Considering 3 to 50 samples of faults for each class the average diagnosis accuracy of the proposed model on TEP is from 81.8% to 94.7%.
@@
Zhang et al. [127]proposed a ZSL approach to handle unseen faults. Zero-shot learning involves the utilization of expert-provided high-level descriptions instead of relying on trained objects to identify target items. These descriptions encompass semantic attributes such as colour, shape, and behaviours, which can be learned beforehand without requiring samples from unfamiliar categories. In essence, ZSL is an approach to train models to recognize patterns in new types of data using information gleaned from familiar classes along with relevant descriptions. At first feature extraction from raw data is done using a one-dimensional convolutional neural network (1D CNN), and then labelling is done by semantic descriptions by expert knowledge. Five Different subsets of TEP are used to validate the proposed method, considering different faults as unseen data. The detection accuracies for different scenarios vary from 59.72% to 96.67%, which outperforms other ZSL methods like SAE in most of the scenarios.
@@
To address the class imbalance problem double branch rebalanced network (DBRN) method is proposed in [54]. At first, a resampling method rebalances the classes, then a cost adaptive reweighting strategy rebalances the cost of each branch. Finally the two aforementioned steps are merged by a fusing learning strategy. The imbalance ratio is 10:1, and for faults 1,2,6 of TEP the average diagnosis accuracy is 98.75%.
@@
Guo et al. [27], used a modified stacked autoencoder to learn low-level features containing the information of complete data based on incomplete input data. This method creates data representations that can be used for classification. The steps of the data driven fault detection process are shown in 11. When the data is not rich, two general approaches can be handle the problem. Augmenting or generating more data, or using the models that can perform with few samples. Different scenarios and possible solutions are shown in the mentioned figure. Table LABEL:tab:Imb shows the summary of methods that can handle imbalanced, incomplete or unlabeled data.
Table 5: Solutions to imbalanced and unlabelled data
Method Reference Specification Fault Information Average Accuracy
@@
Unlabelled [59] Combination of autoencoder and LSTM model and a control chart for anomaly detection of unlabelled data 6.25% of the data is labelled 97.46%
@@
Unlabelled [69] Bidirectional gated recurrent unit (BGRU), weighting training example to overcome the class imbalance and cost-sensitive active learning to explore the unlabelled data Data imbalance ratio of 0.5 and the unlabelled proportion data of 0.6 97.50%
@@
Unlabelled and Imbalanced [67] Enhanced auxiliary classifier generative adversarial network (EACGAN) to deal with imbalanced data and more stable training Normal to fault ratio of 10:1, all faults except 3, 9 and 15 77.01%
@@
Imbalanced [133] Three-step high-fidelity positive-unlabeled (THPU) is used to deal with having a few labels in the data set, by generating and augmenting data One labelled normal sample is available the rest is unlabelled 92.32%
@@
Unlabelled and Imbalanced [65] Contrastive learning for information extraction of unlabeled and imbalanced data Only 20 samples of each fault is available 86.59%
@@
Imbalanced [4] Source-aware autoencoder (SAAE) for handling imbalanced data combining BiLSTM and ResNet Detecting unseen faults 93.80%
@@
Unlabelled [53] Consistency regularization autoencoder (CRAE) for handling unlabelled data All faults except 3, 9, and 15 (20% of data is labelled) 92.95%
@@
Unlabelled and Imbalanced [68] Using genetic algorithm and sparse autoencoder for imbalanced and unlabeled fault diagnosis Unlabelled data and the normal to fault ratio of 10:1 92.02%
@@
Imbalanced [74] High-efficiency GAN using Wasserstein GAN and Auxiliary Classifier, to generate and classify data for imbalanced dataset Normal to fault ratio of 5:1 82.00%
@@
Unlabelled [126] Combined LSTM for temporal feature extraction and ladder autoencoder (LAE) to handle unlabeled data All faults except 3,13, 15 and 16, 11% of the dataset is labelled 95.90%
@@
Unknown Faults [57] Few-shot learning and GRU with attention layer for unknown fault identification Faults 0, 2, 12, and 14 as known faults, Faults 1, 4, 5, 7, 8, 10, 11, and 13 as unknown 85.70%
@@
Imbalanced [101] Few-shot learning based on a class-rebalance strategy to detect new faults. Feature extraction is done using 1D-CNN with an attention module 3 to 50 samples of faults for each class From 81.8% to 94.7%
@@
Imbalanced [54] Double branch rebalanced network (DBRN) to deal with imbalance classes using resampling and reweighting Imbalance ratio is 10:1 for faults 1,2,6 98.75%
5.4 Ensemble Learning
@@
Ensemble learning is a technique that combines the predictions of multiple models to improve the overall performance and robustness of a predictive model. The idea behind ensemble learning is that by aggregating the outputs of several models, you can often achieve better results than any single model can produce [77]. Ensemble methods work by creating a committee of base models, which can be of the same or different types, and then combining their outputs in various ways. Ensemble learning improves the accuracy, robustness and stability of the model. Plakias et al. [71], introduced an unsupervised fault detection approach by the ensemble of different autoencoders, each trained for detecting one type of fault. The final decision is made by a soft voting process. In this approach an anomaly indicator is generated for unknown samples, showing the confidence of the network regarding the condition of the system. To improve the feature learning for complex, nonlinear and high-dimensional signals, a hybrid model of integration of 1-DCNN and stacked denoising auto-encoders (SDAE) is introduced in [119]. This method has higher classification accuracy compared to 1-DCNN, SDAE and DBN. Typically, multiple ML algorithms are combined in ensemble learning; However, as a future direction of research for complex systems like TEP, ensemble learning could be used to aggregate model-based and signal-based methods. This approach could increase the robustness of the condition monitoring systems against uncertainties like imbalanced, incomplete or unlabelled data.
5.5 Explainability

### v3

@@ 4.1.2 Unseen Data
Deep learning-based FDD algorithms can not perform well on unseen operation modes and faults. To deal with previously unknown faults, an algorithm based on supervised contrastive learning (SCL) is used in Peng et al. (2022b), in which the normal samples are contrasted with negative augmentations of themselves. Contrastive Learning is a deep learning technique for unsupervised representation learning. The goal is to learn a representation of data such that similar instances are close together in the representation space, while dissimilar instances are far apart. Soft Brownian Offset (SBO), sampling, and shrinkage autoencoder are used to generate the out-of-distribution (OOD) samples, to obtain some information from unknown classes. The goal is to prevent unseen fault samples from being classified as normal conditions. Balanced fault diagnosis, imbalanced fault diagnosis, and few-shot diagnosis of TEP are tested by the proposed method, demonstrating good performance. Faults 1, 2, 4, 6, and 7 are considered as the seen faults, and faults 8, 12, 14, and 18 are selected as the unseen faults. The detection accuracies in the balanced scenario are 97.5%, 100%, 100%, and 90.63%. In the imbalanced scenario, where the balance ratio of normal to fault is 10:1, the accuracies are 97.2%, 98.75%, 100%, 90.38%. In few-shot learning, where there are only 15 samples of each fault, the accuracies are 95.1%, 97.6%, 99.86%, 88.2%, respectively. Xiao et al. Xiao et al. (2022) implement a developed domain generalization (DG) method that guarantees generalization over unseen domains. DG learns a universal fault diagnosis model from historical operating modes and can generalize to unseen modes. A novel labeling and class progressive adversarial learning (LCPAL) algorithm for FDD is introduced to learn domain-invariant feature representations. This model aligns the label distribution by weighting the source classification loss based on the number of instances in each category. Stacked autoencoders can be used for handling incomplete, missing, or unlabeled data. In Guo et al. (2020), a modified stacked autoencoder is used to learn low-feature representations from incomplete data input. Only 40 input variables out of 52 are randomly chosen for fault detection, and the average accuracy for the fault diagnosis of faults 1, 4, 7, 8, 10, 13 is 83.76%.
A Conditional Gaussian Network (CGN)-based Bayesian Network (BN) framework is proposed in Atoui et al. (2019) for FDD in industrial processes, with the added capability to handle unknown fault classes through a modified exclusion criterion. Unlike traditional approaches that require separate models for detection and diagnosis, the proposed method employs a single BN to simultaneously perform both tasks. By introducing a probabilistic boundary that respects statistical significance levels, the model reduces false alarms and enables statistical fault classification, including the detection of novel faults not seen during training. The approach is validated using TEP under multiple fault scenarios and compared against benchmark methods. Results show improved classification accuracy and a substantial reduction in false alarm rates. Moreover, the proposed framework is computationally simpler and scalable, making it suitable for complex systems with limited labeled fault data.
4.1.3 Imbalanced Data
@@
Industrial systems usually operate under normal condition, hence the normal data is ample. However, faults occur infrequently, and the number of fault samples is rarely equal to the number of normal samples. AI algorithms can not perform well when the data is imbalanced. The performance of the data-driven FDD algorithms depends on the availability of the data; however, in many real-world cases, data is imbalanced or incomplete. To deal with this problem and achieve good FDD accuracy, different approaches are introduced. One approach for handling imbalanced data sets is generating synthetic samples of the minor class to make the data balanced Ahang et al. (2022). The other approach involves using feature extraction methods that focus on available data, as well as considering one-class classification and anomaly detection methods.
@@
Tian et al.Tian et al. (2020), used a generative adversarial network (GAN) to reconstruct missing data, Spearman’s rank correlation coefficient (SRCC) for dimension reduction, select the most important features and noise canceling, and finally DBN for FDD. The average FDR of the process for 20 faults of TEP is 89.70%, which is 6.23% higher than the traditional DBN method. Hu et al.Hu and Jiang (2018) proposed an incremental imbalance modified deep neural network (incremental-IMDNN) in combination with active learning for FDD of imbalanced data. An imbalance modified synthetic minority oversampling technique (IM-SMOTE) is used for addressing data imbalance. The model extends the structure hierarchically with the arrival of new faults. Similar faults are merged into the same classes by using fuzzy clustering. To tackle the class imbalance problem in FDD of chemical processes, an Imbalance Modified Convolutional Neural Network combined with incremental learning, as described in Gu et al. (2021), generates new samples from imbalanced raw data using a dynamic modified synthetic minority oversampling technique (DM-SMOTE). This model updates itself by receiving the new data. In Amini and Zhu (2022), a source-aware autoencoder (SAAE) for detecting unseen faults and handling imbalanced data sets is introduced that uses Bidirectional long short-term memory (BiLSTM) and residual neural network (ResNet) to deal with randomness. The training can be done using just normal data, and then fine-tuned by using emerging faulty samples. The proposed method detects unseen faults with an accuracy of 93.8%; moreover, the average accuracy of fault detection for different fault-to-normal ratios is 91.8%. In the balanced condition, the average accuracy is 94.3%. Peng et al. Peng et al. (2022c) introduced a non-revisiting genetic cost-sensitive sparse autoencoder(NrGCS-SAE) for solving the imbalanced fault diagnosis problem. The genetic algorithm optimizes class weights while the non-revisiting strategy increases exploration ability. This method could detect faults in the unbalanced TE dataset with an average accuracy of 92.02% while the normal-to-fault ratio was 10:1.
@@
To identify unknown faults, a few-shot learning-based unknown recognition and classification (FSLB-UR&C) approach is introduced in Mirzaei et al. (2023). For feature extraction and comparing the similarity of features, a gated recurrent unit with a dot-product attention layer is used. This method does not require a large amount of data for training. The classification accuracy of the known faults of TEP is 88.9%, and for unknown faults is 85.7%. The known faults include Faults 0, 2, 12, and 14, while the unknown faults include Faults 1, 4, 5, 7, 8, 10, 11, and 13. Xu et al. Xu et al. (2023), propose a few-shot learning fault diagnosis method based on a class-rebalance strategy, to diagnose new scarce faults. The fault samples are transferred into a feature space, where each type of fault, even each new instance, is mapped into a separate feature cluster. In this way, fault diagnosis can be achieved by estimating feature similarity between instances and faults. 1D-CNN with an attention module is used for feature extraction. Data augmentation and class rebalance are done to increase the fault samples of scarce classes. Considering 3 to 50 samples of faults for each class the average diagnosis accuracy of the proposed model on TEP is from 81.8% to 94.7%.
@@
Zhang et al. Zhang et al. (2023)proposed a ZSL approach to handle unseen faults. Zero-shot learning involves the utilization of expert-provided high-level descriptions instead of relying on trained objects to identify target items. These descriptions encompass semantic attributes such as colour, shape, and behaviors, which can be learned beforehand without requiring samples from unfamiliar categories. In essence, ZSL is an approach to train models to recognize patterns in new types of data using information gleaned from familiar classes along with relevant descriptions. Initially, feature extraction from raw data is performed using a one-dimensional convolutional neural network (1D CNN), followed by labeling with semantic descriptions based on expert knowledge. Five Different subsets of TEP are used to validate the proposed method, considering different faults as unseen data. The detection accuracies for different scenarios vary from 59.72% to 96.67%, which outperforms other ZSL methods like SAE in most of the scenarios.
@@
To address the class imbalance problem, the double branch rebalanced network (DBRN) method is proposed in Ma et al. (2023). At first, a resampling method rebalances the classes, then a cost adaptive reweighting strategy rebalances the cost of each branch. Finally, the two aforementioned steps are merged by a fusing learning strategy. The imbalance ratio is 10:1, and for faults 1,2,6 of TEP the average diagnosis accuracy is 98.75%.
@@
Guo et al. Guo et al. (2020), used a modified stacked autoencoder to learn low-level features containing the information of complete data based on incomplete input data. This method creates data representations that can be used for classification. The strategies for the data-driven fault detection process are shown in 5. When the data is not rich, two general approaches can handle the problem. Augmenting or generating more data, or using the models that can perform with few samples. Different scenarios and possible solutions are shown in the mentioned figure. Table 8 shows the summary of methods that can handle imbalanced, incomplete, or unlabeled data.
Table 8: Solutions to imbalanced and unlabeled data
Method Reference Specification Fault Information Avg. Accuracy
@@
unlabeled Morales-Forero and Bassetto (2019) Combination of autoencoder and LSTM model and a control chart for anomaly detection of unlabeled data 6.25% of the data is labeled 97.46%
@@
unlabeled Peng et al. (2020b) Bidirectional gated recurrent unit (BGRU), weighting training example to overcome the class imbalance and cost-sensitive active learning to explore the unlabeled data Data imbalance ratio of 0.5 and the unlabeled proportion data of 0.6 97.50%
@@
unlabeled and Imbalanced Peng et al. (2020a) Enhanced auxiliary classifier generative adversarial network (EACGAN) to deal with imbalanced data and more stable training Normal to fault ratio of 10:1, all faults except 3, 9 and 15 77.01%
@@
Imbalanced Zheng and Zhao (2022) Three-step high-fidelity positive-unlabeled (THPU) is used to deal with having a few labels in the data set, by generating and augmenting data One labeled normal sample is available the rest is unlabeled 92.32%
@@
unlabeled and Imbalanced Peng et al. (2022a) Contrastive learning for information extraction of unlabeled and imbalanced data Only 20 samples of each fault is available 86.59%
@@
Imbalanced Amini and Zhu (2022) Source-aware autoencoder (SAAE) for handling imbalanced data combining BiLSTM and ResNet Detecting unseen faults 93.80%
@@
unlabeled Ma et al. (2022) Consistency regularization autoencoder (CRAE) for handling unlabeled data All faults except 3, 9, and 15 (20% of data is labeled) 92.95%
@@
unlabeled and Imbalanced Peng et al. (2022c) Using genetic algorithm and sparse autoencoder for imbalanced and unlabeled fault diagnosis unlabeled data and the normal to fault ratio of 10:1 92.02%
@@
Imbalanced Qin and Zhao (2022b) High-efficiency GAN using Wasserstein GAN and Auxiliary Classifier, to generate and classify data for an imbalanced dataset Normal to fault ratio of 5:1 82.00%
@@
unlabeled Zhang and Qiu (2022b) Combined LSTM for temporal feature extraction and ladder autoencoder (LAE) to handle unlabeled data All faults except 3,13, 15, and 16, 11% of the dataset is labeled 95.90%
@@
Unknown Faults Mirzaei et al. (2023) Few-shot learning and GRU with attention layer for unknown fault identification Faults 0, 2, 12, and 14 as known faults, Faults 1, 4, 5, 7, 8, 10, 11, and 13 as unknown 85.70%
@@
Imbalanced Xu et al. (2023) Few-shot learning based on a class-rebalance strategy to detect new faults. Feature extraction is done using 1D-CNN with an attention module 3 to 50 samples of faults for each class From 81.8% to 94.7%
@@
Imbalanced Ma et al. (2023) Double branch rebalanced network (DBRN) to deal with imbalanced classes using resampling and reweighting Imbalance ratio is 10:1 for faults 1,2,6 98.75%
4.2 Emerging Methods and Future Directions
This section highlights emerging methodologies and concepts that have significant potential for improving intelligent condition monitoring. While these approaches have demonstrated promising results, they remain relatively underexplored in industrial applications. By examining these cutting-edge approaches, we aim to identify future research opportunities that can address current limitations, leading to the development of more reliable, interpretable, and scalable condition monitoring systems.
4.2.1 Ensemble Learning
@@
Ensemble learning is a technique that combines the predictions of multiple models to improve the overall performance and robustness of a predictive model. The idea behind ensemble learning is that by aggregating the outputs of several models, you can often achieve better results than any single model can produce Sagi and Rokach (2018). Ensemble methods work by creating a committee of base models, which can be of the same or different types, and then combining their outputs in various ways. Ensemble learning improves the accuracy, robustness, and stability of the model. Plakias et al. Plakias and Boutalis (2022) introduced an unsupervised fault detection approach by the ensemble of different autoencoders, each trained for detecting one type of fault. A soft voting process makes the final decision. In this approach, an anomaly indicator is generated for unknown samples, showing the confidence of the network regarding the condition of the system. To improve the feature learning for complex, nonlinear, and high-dimensional signals, a hybrid model of integration of 1-DCNN and stacked denoising auto-encoders (SDAE) is introduced in Zhang et al. (2021). This method has higher classification accuracy compared to 1-DCNN, SDAE, and DBN.
Li et al. Li et al. (2022) proposed a novel ensemble monitoring framework for industrial process data to address the complexities inherent in industrial processes, which exhibit various characteristics such as linearity, non-linearity, Gaussian, non-Gaussian, and dynamic behavior. The framework aims to automatically determine the most suitable local models and optimal monitoring variables by using a combination of multiple statistical analysis methods. The key contributions of the study include the automatic determination of local models and monitoring variables, the use of multiple models to capture diverse data characteristics, and the elimination of redundant models to enhance monitoring performance. An ensemble convolutional neural network is introduced in Najaran (2023), which combines base learner algorithms trained on features extracted by different approaches such as Fourier transforms, Wavelet transforms, Walsh transforms, and Hilbert–Huang transforms. CNNs can also automatically extract features from these signals. The output of base learners is used in a weighted voting paradigm, which is optimized via an evolutionary algorithm. Typically, multiple ML algorithms are combined in ensemble learning; However, as a future direction of research for complex systems like TEP, ensemble learning could be used to aggregate model-based and signal-based methods. This approach could increase the robustness of the condition monitoring systems against uncertainties like imbalanced, incomplete, or unlabeled data.
4.2.2 Hybrid and Informed Machine Learning
Data-driven models highly depend on the availability and quality of the data, which is very difficult to acquire in complex industrial setups such as chemical plants. By integrating prior knowledge into data-driven and ML methods, and utilizing hybrid and informed approaches, we can overcome the data scarcity and enhance the accuracy and safety of the process Wu et al. (2024); Wilhelm et al. (2021). Hybrid FDD approaches that combine model-based reasoning with data-driven learning are gaining attention in the context of safety-critical industrial systems. These methods leverage both the interpretability of physical or probabilistic models and the adaptability of machine learning techniques. Informed machine learning describes learning from a hybrid information source that consists of data and prior knowledge. The prior knowledge can be integrated into machine learning in several stages, including training data, hypothesis set, learning algorithm, or final decision Von Rueden et al. (2021).
Alauddin et al. Alauddin et al. (2023) integrate process dynamics into data-driven models to improve the model generalization when the available data is sparse. An additional layer is added to the neural network architecture to incorporate equations and field knowledge, introducing a process dynamics-guided neural network (PDNN) model. Compared to purely data-driven neural networks, PDNN shows improvement on reduced sample-sized data. However, an accurate model of the process dynamics is needed to achieve acceptable performance, and such models are computationally more expensive. Chao et al. Chao et al. (2022) introduce a hybrid framework using physics-based models to infer unobservable model parameters, which are later combined with sensor readings and used as an input to a deep neural network for prognostics of complex systems. The framework was implemented for turbofan engine remaining useful lifetime (RUL) prediction, and outperformed purely data-driven approaches and required less training data. Tidriri et al. Tidriri et al. (2018) introduced a framework for decision fusion of heterogeneous methods for FDD to obtain reliable results. The proposed discrete Bayesian Network (BN) approach is validated using TEP, combining model-based and data-driven approaches, and outperforms the individual decisions, achieving an FDR of 99.21%.
4.2.3 Explainability